# The shrinking target and recurrence problem for non-autonomous systems

### v1

@@ Abstract.
We investigate the shrinking target and recurrence set associated to non-autonomous measure-preserving systems on compact metric spaces, establishing zero–one criteria in the spirit of classical Borel–Cantelli results.
Our first main theorem gives a quantitative shrinking target result for non-autonomous systems under a uniform mixing condition, providing asymptotics with an optimal error term. This general result is applicable to certain families of inner functions, yielding concrete applications such as digit patterns in multi-base expansions. Turning to recurrence, we establish new zero-measure laws for non-autonomous systems. In the autonomous case, we prove a zero–one criterion for recurrence sets of centred, one-component inner functions via Markov partitions and distortion estimates. Together, these results provide a unified framework for shrinking target and recurrence problems in both autonomous and non-autonomous dynamics.
@@ 1. Introduction
In contrast, study of the non-autonomous shrinking target problem is far less complete. Even when $r_{n}=c$, and $t_{n}$ ergodic for all $n$, full measure of $\mathcal{W}(T_{\mathbb{N}},\{x_{n},r_{n}\})$ and $\mathcal{R}(T_{\mathbb{N}},\{r_{n}\})$ need not hold, since compositions of measure-preserving maps can degrade ergodic properties–see Section 5 for discussion. Nevertheless, recent work has begun to address such problems. In [12], Fishman-Mance-Simmons-Urbański study the Cantor series expansion of points in $[0,1]$, which we define later before stating Theorem C. They give bounds on the Hausdorff dimension of the associated non-autonomous shrinking target set. Subsequently, in the same context, Cao-Sun [29, Theorem 1.1] provides a zero-one criteria for the set .
In complex dynamics, attention has turned to shrinking target problems for non-autonomous systems of inner functions. These functions are holomorphic self-maps of the unit disk that extend Lebesgue almost everywhere to maps of the boundary (see Section 2.3, Definition 2.4), and are closed under composition [27, Theorem 2]. They provide a natural framework for bounded holomorphic dynamics as every bounded holomorphic map admits a canonical outer–inner function factorisation. A further motivation of their study comes from investigating entire maps with simply connected wandering domains of entire functions. Here, under iteration, the domain ‘wanders’ through $\mathbb{C}$}, and repeated application of the Riemann Mapping Theorem transfers the dynamics to the closed unit disk, inducing a non-autonomous system of inner functions. This perspective underlies works such as [3], in which the authors establish a zero-one criterion as well as a variety of zero and full measure laws in this non-autonomous inner function setting. Here, we use the term zero (or full) measure law to refer to a result that provides sufficient conditions under which the set of interest has zero (or full) measure.
@@
Setting $t_{n}=t$ and $x_{n}=x_{0}$ for all $n$, we recover the notion of uniform mixing at a point introduced in [9, Definition 2]. If we set $\phi(n)=\gamma^{n}$, for $0<\gamma<1$, then $T_{\mathbb{N}}$ is exponentially mixing in the usual sense.
@@
Theorem A is a quantitative version of [9, Theorem 1], and a generalisation of [22, Theorem 1] to the non-autonomous setting. As a concrete application, we specialise Theorem A to a system of inner functions. Throughout, $\mathbb{D}$} will denote the open unit disk in $\mathbb{C}$}, $\partial{\mathbb{D}}$ its boundary (i.e. the unit circle) and $\lambda$ is normalised Lebesgue measure. Consider a sequence of centred, uniformly contracting inner functions $F_{\mathbb{N}}:=\{f_{n}\}_{n\in{\mathbb{N}}}$, meaning $f_{n}(0)=0$ and $|f_{n}^{\prime}(0)|\leq\alpha<1$ for all $n\in{\mathbb{N}}$. Then, we will show in Section 2.3 that $(\partial{\mathbb{D}},\mathcal{B},\lambda,F_{\mathbb{N}},d)$ is an exponentially mixing system, where $d$ normalised arc length.
@@
As an application of Theorem B, we prove a result describing the occurrence of an arbitrarily long prescribed string of digits in the multi-base or ‘Cantor series’ expansion of points in the unit interval, first introduced by Cantor in [5]. This expansion is the unique expansion of $x\in[0,1)$ with respect to the sequence of bases $\{b_{i}\}_{i\in{\mathbb{N}}}$ with $b_{i}\in{\mathbb{N}}_{\geq 2}$ such that
@@
where the digits $d_{i}\neq b_{i}-1$ for infinitely many $i\in{\mathbb{N}}$. We will denote this expansion by $x=[d_{1},..,d_{n},\dots]$ and with this notation in place, we can state our theorem.
@@
Let $\{b_{i}\}_{i\in{\mathbb{N}}}$ be a sequence of bases with $b_{i}\in{\mathbb{N}}_{\geq 2}$, and $\xi:{\mathbb{N}}_{>0}\to{\mathbb{N}}_{>0}$ be an increasing function. Then, for any given $\varepsilon>0$, we have that for almost every $x=[d_{1},\dots,d_{n},\dots]\in[0,1)$,
@@
where $\Phi(N):=\sum\limits_{n=1}^{N}\frac{1}{b_{1}\dotsb_{\xi(n)+1}}$. In particular,
@@
where $D(\xi)$ is the set of $x$ for which $d_{n+1}=d_{n+2}=\dots=d_{n+\xi(n)}=0$ for infinitely many $n\in{\mathbb{N}}$
We actually obtain a stronger version of Theorem C. Indeed, given a real number $x_{0}$, consider its multi-base expansion $x_{0}=[c_{1},\dots,c_{n},\dots]$ with respect to the sequence of bases $\{b_{i}\}_{n\in{\mathbb{N}}}$. Then we have the same statement as in Theorem C, but with $d_{n+1}=c_{1},\dots,d_{n+\xi(n)}=c_{\xi(n)}$. This result is given in full as Theorem G in Section 2.6. It is both a generalisation and strengthening of [9, Theorem 5] which considers expansions with respect to a fixed base $b$, and also of [29, Theorem 1.1].
We now turn our attention to the recurrent set $\mathcal{R}(T_{\mathbb{N}},\{r_{n}\})$. Recall that for (measure-preserving) autonomous systems $(X,\mathcal{B},\mu,t,d)$, Poincaré Recurrence guarantees $\mu(\mathcal{R}(t,\{r_{n}\}))=1$ when $\{r_{n}\}_{n\in{\mathbb{N}}}$ is a constant sequence. In the case $r_{n}\to 0$, in pioneering work, Boshernitzan [4] established full measure results for $\mathcal{R}(t,\{r_{n}\})$ under mild additional assumptions on $X$, beyond mere measure-preservation. Complete zero–one criteria, however, have only been obtained for autonomous dynamical systems that, at minimum, are uniformly mixing (see Definition 1.3) and admit a ‘regular’ partition (see Section 4.2). Such systems include those considered within [1], [6], [15], [21] and, particularly relevant to this paper, [16] in which He-Liao considered piecewise expanding maps with a finite partition, and [18], in which Hussain-Li-Simmons-Wang consider maps with a countable partition and regularity properties. The study of non-autonomous recurrence, however, remains virtually unexplored. We remedy this gap with the following result, which provides a zero measure criteria for the recurrence set of a non-autonomous system.
@@
This proof is split into two cases, depending on if $f$ is of finite or infinite degree. The proof in the finite degree case follows an application of the previously mentioned [16, Theorem 1.3]. The proof for infinite degree one-component inner functions relies on the fact that they admit a countable Markov partition as shown in [19]. We recall this construction in Section 4.2, and we then prove that the partition satisfies various distortion estimates necessary to prove the theorem. The overall strategy follows that of [18, Theorem 1.3], but since the extension of $f$ to the unit circle fails to be a self-map on a set of Lebesgue measure zero, the result does not follow directly, and requires non-trivial work.
@@ 2.1. Borel-Cantelli Lemmas
The proof hinges on the Cauchy-Schwarz Inequality, see [14, Lemma 3] for details. We refer to (2.1) as a ‘quasi-independence on average’ condition. Lemma 2.2 only allows us to conclude positive, not full measure, unless $C=1$. However, often we can jump from positive measure to a full measure result via theorems such as Kolmorogorov’s Zero-One Law.
The following lemma, which can be found in [13, pg 13], is sometimes referred to as a quantitative Borel-Cantelli Lemma.
@@ 2.3. Background on inner functions
From Lemma 2.6 and Lemma 2.7, we conclude that if $f$ is centred, then $(\partial{\mathbb{D}},\mathcal{B},\lambda,f,d)$ is a uniformly expanding, $\lambda$-measure-preserving dynamical system on $\partial{\mathbb{D}}$.
We say a centred inner function is contracting (in $\mathbb{D}$}) if $|f^{\prime}(0)|=\alpha<1$. Given a sequence $F_{\mathbb{N}}$ of centred inner function, we say $F_{\mathbb{N}}$ is contracting if $|f_{n}^{\prime}(0)|=\alpha_{n}<1$ for all $n\in{\mathbb{N}}$, and, as mentioned in the introduction, uniformly contracting if $|f_{n}^{\prime}(0)|\leq\alpha<1$. For centred inner functions, $\alpha_{n}$ is the hyperbolic distortion of $f$ at 0. More generally, we say a sequence of inner functions is contracting if $\sum_{n=1}^{\infty}(1-\alpha_{n})=\infty$, in which case the hyperbolic distortion $\alpha_{n}$ may have a more complicated expression– see [3] and the references within for details. Therefore, centred inner functions always act as contractions on $\mathbb{D}$} and as expansions on $\partial{\mathbb{D}}$; moreover, uniform contraction on $\mathbb{D}$} is equivalent to uniform expansion on $\partial{\mathbb{D}}$.
@@
The uniform contracting condition $|f_{n}^{\prime}(0)|\leq\alpha<1$ ensures that the compositional map $F_{n}$ satisfies, $|F_{n}^{\prime}(0)|\leq\alpha^{n}$ for each $n\in{\mathbb{N}}$. In [3], Benini-Evdoridou-Fagella-Rippon-Stallard obtain full measure shrinking target results for compositions of inner functions with control on the hyperbolic distortion under the assumption that a sum involving the terms $r_{n}$ and $\alpha_{n}$ diverges.
@@ 2.5. Background on base expansions
We provide a short review on base expansions, needed to understand the statement and proof of Theorem C. We start by considering maps of the form
@@
Notice that our condition on $2\leq b\in{\mathbb{N}}$ is essential to establish the link to symbolic dynamics. We give an example to indicate that this result does not hold for all $b\in{\mathbb{R}}$.
@@
This isomorphism is used to translate between the study of inner functions and base expansions– see, for example, [9, Theorem 4.5].
@@
As described in the introduction, the multi-base or ‘Cantor series’ expansion of $x$ with respect to $\{b_{i}\}_{n\in{\mathbb{N}}}$, $b_{i}\in{\mathbb{N}}_{\geq 2}$ is the unique expansion of $x\in[0,1)$:
@@
where $d_{i}\neq b_{i}-1$ for infinitely many $i$, and is denoted by $x=[d_{1},..,d_{i},\dots]$. In a similar way as to for a fixed base $b$, we can view $T_{n}$ as a left shift map on the expansion of $x$ with respect to $\{b_{i}\}_{i\in{\mathbb{N}}}$:
$T_{1}([d_{1},\dots,d_{i},\dots])\,=\,[d_{2},\dots,d_{i+1},\dots]$
This concludes the background material on base expansion.
@@ 2.6. Proof of Theorem C
Theorem C follows from a non trivial application of Theorem B. As discussed in the introduction, we shall actually prove a more general version of Theorem C. We state this now, using the notation established in Section 2.5 for the Cantor series expansion of points in the unit interval.
Theorem G.
Let $\{b_{i}\}_{i\in{\mathbb{N}}}$ be a sequence of bases with $b_{i}\in{\mathbb{N}}_{\geq 2}$, and $\xi:{\mathbb{N}}\to{\mathbb{N}}$ be an increasing function. Then, given a point $x_{0}=[c_{1},\dots,c_{i},\dots]\in[0,1)$, for any $\epsilon>0$ we have
$\big\{1\leq n\leq N:d_{n+1}=c_{1},\dots,d_{n+\xi(n)}=c_{\xi(n)}\big\}=\Phi(N)+O\left(\Phi^{1/2}(N)\ (\log\Phi(N))^{3/2+\varepsilon}\right)$
for almost every $x=[d_{1},\dots,d_{i},\dots]$, where $\Phi(N):=\sum\limits_{n=1}^{N}\frac{1}{b_{1}\dotsb_{\xi(n)+1}}$. In particular,
[…]
where $D_{x_{0}}(\xi)$ is the set of $x$ for which $d_{n+1}=c_{1},\dots,d_{n+\xi(n)}=c_{\xi(n)}$ for infinitely many $n\in{\mathbb{N}}$.
Proof of Theorem G.
@@
We now illustrate the link between the shrinking target set and the digit set $D_{x_{0}}(\xi)$. Given a point $x=[d_{1},\dots,d_{n},\dots]\in[0,1)$, the map $T_{n}$ acts as a shift map on $x$:
$T_{n}(x)=[d_{n+1},\dots]=\frac{d_{n+1}}{b_{1}}\ +\ \frac{d_{n+2}}{b_{1}b_{2}}\ +\ \dots\ +\ \frac{d_{n+\xi(n)}}{b_{1}\dotsb_{\xi(n)}}\ +\ \frac{d_{n+\xi(n)+1}}{b_{1}\dotsb_{\xi(n)+1}}\ +\ \dots\,$
To force $d_{n+1}=c_{1},d_{n+2}=c_{2},\dots,d_{n+\xi(n)}=c_{\xi(n)}$, we require $|T_{n}(x)-x_{0}|\leq\frac{1}{b_{1}\dotsb_{\xi(n)+1}}$. We can interpret this condition through the shrinking target framework as
$|T_{n}(x)-x_{0}|\ \leq\ \textstyle{\frac{1}{b_{1}\dotsb_{\xi(n)+1}}}\quad\iff\quad T_{n}(x)\ \in\ B\left(x_{0},\frac{1}{b_{1}\dotsb_{\xi(n)+1}}\right)\,.$
Therefore, with respect to (1.1), pick $x_{n}=x_{0}$ for all $n$ and $r_{n}=1/(b_{1}\dotsb_{\xi(n)+1})$. On applying Theorem B to the shrinking target set $\mathcal{W}(T_{\mathbb{N}},\{x_{0},r_{n}\})$, then we are done. ∎
With reference to Theorem G, note that
${\Phi(N)=\sum_{n=1}^{N}\frac{1}{b_{1}\dotsb_{\xi(n)+1}}\leq\sum_{n=1}^{N}2^{-(\xi(n)+1)}}\,$,
and so, on choosing, for example, $\xi(n)=n$, the sum to infinity converges. Conversely, on choosing, for example, $\xi(n)=b_{1}\cdots b_{l}$ where $l\in{\mathbb{N}}$ is maximum integer for which $b_{1}\cdots b_{l}<n$, the sum to infinity diverges.
Finally, we obtain Theorem C from setting $x_{0}=0$ in Theorem G.

### v2

@@ Abstract.
We investigate the shrinking target and recurrence set associated to non-autonomous measure-preserving systems on compact metric spaces, establishing zero-one criteria in the spirit of classical Borel-Cantelli results.
Our first main theorem gives a quantitative shrinking target result for non-autonomous systems under a uniform mixing condition, providing asymptotics with an optimal error term. This general result is applicable to certain families of inner functions, yielding concrete applications such as patterns of zeros in the multibase expansion. Turning to recurrence, we establish new zero-measure laws for non-autonomous systems. In the autonomous case, we prove a zero-one criterion for recurrence sets of centred, one-component inner functions via Markov partitions and distortion estimates. Together, these results provide a unified framework for shrinking target and recurrence problems in both autonomous and non-autonomous dynamics.
@@ 1. Introduction
In contrast, study of the non-autonomous shrinking target problem is far less complete. Even when $r_{n}=c$, and $t_{n}$ ergodic for all $n$, full measure of $\mathcal{W}(T_{\mathbb{N}},\{x_{n},r_{n}\})$ and $\mathcal{R}(T_{\mathbb{N}},\{r_{n}\})$ need not hold, since compositions of measure-preserving maps can degrade ergodic properties–see Section 5 for discussion. Nevertheless, recent work has begun to address such problems. In [11], Fishman-Mance-Simmons-Urbański study the Cantor series expansion of points in $[0,1]$, which we define later before stating Theorem C. They give bounds on the Hausdorff dimension of the associated non-autonomous shrinking target set. Subsequently, in the same context, Sun [29, Theorem 1.1] provides a zero-one criteria for the set.
In complex dynamics, attention has turned to shrinking target problems for non-autonomous systems of inner functions. These functions are holomorphic self-maps of the unit disk that extend Lebesgue almost everywhere to maps of the boundary (see Section 2.3, Definition 2.4), and are closed under composition [27, Theorem 2]. They provide a natural framework for bounded holomorphic dynamics as every bounded holomorphic map admits a canonical outer–inner function factorisation. A further motivation of their study comes from investigating entire maps with simply connected wandering domains of entire functions. Here, under iteration, the domain ‘wanders’ through $\hat{\mathbb{C}}$, and repeated application of the Riemann Mapping Theorem transfers the dynamics to the closed unit disk, inducing a non-autonomous system of inner functions. This perspective underlies works such as [3], in which the authors establish a zero-one criterion as well as a variety of zero and full measure laws in this non-autonomous inner function setting. Here, we use the term zero (or full) measure law to refer to a result that provides sufficient conditions under which the set of interest has zero (or full) measure.
@@
Setting $t_{n}=t$ and $x_{n}=x_{0}$ for all $n$, we recover the notion of uniform mixing at a point introduced in [8, Definition 2]. If we set $\phi(n)=\gamma^{n}$, for $0<\gamma<1$, then $T_{\mathbb{N}}$ is exponentially mixing in the usual sense.
@@
Theorem A is a quantitative version of [8, Theorem 1], and a generalisation of [21, Theorem 1] to the non-autonomous setting. As a concrete application, we specialise Theorem A to a system of inner functions. Throughout, $\mathbb{D}$} will denote the open unit disk in $\mathbb{C}$}, $\partial{\mathbb{D}}$ its boundary (i.e. the unit circle) and $\lambda$ is normalised Lebesgue measure. Consider a sequence of centred, uniformly contracting inner functions $F_{\mathbb{N}}:=\{f_{n}\}_{n\in{\mathbb{N}}}$, meaning $f_{n}(0)=0$ and $|f_{n}^{\prime}(0)|\leq\alpha<1$ for all $n\in{\mathbb{N}}$. Then, we will show in Section 2.3 that $(\partial{\mathbb{D}},\mathcal{B},\lambda,F_{\mathbb{N}},d)$ is an exponentially mixing system, where $d$ is normalised arc length.
@@
As an application of Theorem B, we prove a zero-one criterion on the occurrence of arbitrarily long string of zeros in the multibase or ‘Cantor series’ expansion of points in the unit interval. This expansion is the unique expansion of $x\in[0,1)$ with respect to the sequence of bases $\{b_{i}\}_{i\in{\mathbb{N}}}$ with $b_{i}\in{\mathbb{N}}_{\geq 2}$ such that
@@
where the digits $d_{i}\neq b_{i}-1$ for all $i\geq i_{0}\in{\mathbb{N}}$. We will denote this expansion by $x=[d_{1},..,d_{n},\dots]$ and with this notation in place, we can state our theorem.
@@
Let $\{b_{i}\}_{i\in{\mathbb{N}}}$ be a sequence of bases with $b_{i}\in{\mathbb{N}}_{\geq 2}$, and $\xi:{\mathbb{N}}\to{\mathbb{N}}$ be an increasing function. Then, for any given $\varepsilon>0$, we have that for almost every $x=[d_{1},\dots,d_{n},\dots]\in[0,1)$,
@@
where $\Phi(N):=\sum\limits_{n=1}^{N}\frac{1}{b_{n+1}\dots b_{n+1+\xi(n)}}$. In particular,
@@
where $D(\xi)$ is $x=[d_{1},\dots,d_{n}]\in[0,1)$ for which $d_{n+1}=d_{n+2}=\dots=d_{n+\xi(n)}=0$ for infinitely many $n\in{\mathbb{N}}$.
In [23], Mance provides an asymptotic result pertaining to the typicality of normal numbers in the multibase expansion for sequence of bases tending to infinity. More precisely, he quantifies the frequency of blocks (not necessarily zeros) of fixed length in such expansions. Therefore, the more ‘interesting’ case of Theorem C is when the function $\xi$ is unbounded, since, in this case, the result concerns the occurrence of arbitrarily long sequences of zeros in the multibase expansion. We discuss this further following the proof of Theorem C in Section 2.6. Note that on specialising to $b_{i}=b$ for all $i$, we return to the autonomous case, where our result concerns the base $b$ expansion of points $x\in[0,1)$. In this case, we can readily strengthen the conclusion of [8, Theorem 5], a related result concerning the frequencies of arbitrarily long strings of digits (not necessarily zeros) in the base $b$ expansion, by upgrading it to an asymptotic statement with optimal error.
We now turn our attention to the recurrence set $\mathcal{R}(T_{\mathbb{N}},\{r_{n}\})$. Recall that for (measure-preserving) autonomous systems $(X,\mathcal{B},\mu,t,d)$, Poincaré Recurrence guarantees $\mu(\mathcal{R}(t,\{r_{n}\}))=1$ when $\{r_{n}\}_{n\in{\mathbb{N}}}$ is a constant sequence. In the case $r_{n}\to 0$, in pioneering work, Boshernitzan [4] established full measure results for $\mathcal{R}(t,\{r_{n}\})$ under mild additional assumptions on $X$, beyond mere measure-preservation. Complete zero–one criteria, however, have only been obtained for autonomous dynamical systems that, at minimum, are uniformly mixing (see Definition 1.3) and admit a ‘regular’ partition (see Section 4.2). Such systems include those considered within [1], [5], [14], [20] and, particularly relevant to this paper, [15] in which He-Liao considered piecewise expanding maps with a finite partition, and [17], in which Hussain-Li-Simmons-Wang consider maps with a countable partition and regularity properties. The study of non-autonomous recurrence, however, remains virtually unexplored. We remedy this gap with the following result, which provides a zero measure criteria for the recurrence set of a non-autonomous system.
@@
This proof is split into two cases, depending on if $f$ is of finite or infinite degree. The proof in the finite degree case follows an application of the previously mentioned [15, Theorem 1.3]. The proof for infinite degree one-component inner functions relies on the fact that they admit a countable Markov partition as shown in [18]. We recall this construction in Section 4.2, and we then prove that the partition satisfies various distortion estimates necessary to prove the theorem. The overall strategy follows that of [17, Theorem 1.3], but since the extension of $f$ to the unit circle fails to be a self-map on a set of Lebesgue measure zero, the result does not follow directly, and requires non-trivial work.
@@ 2.1. Borel-Cantelli Lemmas
The proof hinges on the Cauchy-Schwarz Inequality, see [13, Lemma 3] for details. We refer to (2.1) as a ‘quasi-independence on average’ condition. Lemma 2.2 only allows us to conclude positive, not full measure, unless $C=1$. However, often we can jump from positive measure to a full measure result via theorems such as Kolmorogorov’s Zero-One Law.
The following lemma, which can be found in [12, pg 13], is sometimes referred to as a quantitative Borel-Cantelli Lemma.
@@ 2.3. Background on inner functions
From Lemma 2.6 and Lemma 2.7, we conclude that if $f$ is centred, then $(\partial{\mathbb{D}},\mathcal{B},\lambda,f,d)$ is a uniformly expanding, $\lambda$-measure-preserving dynamical system on $\partial{\mathbb{D}}$. We say a centred inner function is contracting (in $\mathbb{D}$}) if $|f^{\prime}(0)|=\alpha<1$. Given a sequence $F_{\mathbb{N}}$ of centred inner function, we say $F_{\mathbb{N}}$ is contracting if $|f_{n}^{\prime}(0)|=\alpha_{n}<1$ for all $n\in{\mathbb{N}}$, and, as mentioned in the introduction, uniformly contracting if $|f_{n}^{\prime}(0)|\leq\alpha<1$. For centred inner functions, $\alpha_{n}$ is the hyperbolic distortion of $f$ at 0. More generally, we say a sequence of inner functions is contracting if $\sum_{n=1}^{\infty}(1-\alpha_{n})=\infty$, in which case the hyperbolic distortion $\alpha_{n}$ may have a more complicated expression– see [3] and the references within for details. Therefore, centred inner functions always act as contractions on $\mathbb{D}$} and as expansions on $\partial{\mathbb{D}}$; moreover, uniform contraction on $\mathbb{D}$} is equivalent to uniform expansion on $\partial{\mathbb{D}}$.
@@
The uniform contracting condition $|f_{n}^{\prime}(0)|\leq\alpha<1$ ensures that the compositional map $F_{n}$ satisfies, $|F_{n}^{\prime}(0)|\leq\alpha^{n}$ for each $n\in{\mathbb{N}}$. In [3], Benini-Evdoridou-Fagella-Rippon-Stallard study inner functions which are contracting (that is $|f_{n}^{\prime}(0)|=\alpha_{n}<1$ for each $n\in{\mathbb{N}}$) rather than uniformly contracting, and show that a full measure shrinking target result follows if the sum of $r_{n}(1-\alpha_{n})$ diverges, where $r_{n}$ is the sequence of radii in (1.1).
@@ 2.5. Background on base expansions
With Theorem C in mind, we provide an appropriate overview of base expansions. We start by considering maps of the form
@@
Notice that our condition on $2\leq b\in{\mathbb{N}}$ is essential to establish the link to symbolic dynamics. We give an example to indicate that this “link” does not hold for all $b\in{\mathbb{R}}_{\geq 2}$.
@@
This isomorphism is used to translate between the study of inner functions and base expansions– see, for example, [8, Theorem 4.5].
@@
As described in the introduction, the multibase or ‘Cantor series’ expansion of $x\in[0,1)$ with respect to $\{b_{i}\}_{n\in{\mathbb{N}}}$, $b_{i}\in{\mathbb{N}}_{\geq 2}$ is the unique expansion:
@@
where $d_{i}\neq b_{i}-1$ for all $i\geq i_{0}$, and is denoted by $x=[d_{1},..,d_{i},\dots]$. In contrast to the fixed base case $b$, the transformation $T_{n}$ can no longer be interpreted as a simple left-shift of the digit string; that is to say:
$T_{1}([d_{1},\dots,d_{i},\dots])\,\neq\,[d_{2},\dots,d_{i+1},\dots]\,.$
Indeed, in the multibase setting, we have
$T_{n}([d_{1},\ldots,d_{n},\ldots])=\frac{d_{n+1}}{b_{n+1}}+\frac{d_{n+2}}{b_{n+1}b_{n+2}}+\cdots\,.$
Therefore, the map $T_{n}$ acts as a left shift not only on the digits but also on the sequence of bases itself; it replaces the origional base sequence $\{b_{i}\}_{i\in{\mathbb{N}}}$ by the shifted sequence $\{b_{n+i}\}_{i\in{\mathbb{N}}}$. As in the autonomous setting, the sequence of bases was the constant sequence $b_{n}=b$ for all $n\in{\mathbb{N}}$, this shift in the bases is invisible; there is no effective change in the underlying expansion mechanism. However, as the proof of Theorem C reveals, in the non-autonomous setting the shift of the base sequence is exactly what restricts the result to sequences of zeros, rather than to general digit sequences.
Finally, we conclude the section by mentioning [29, Theorem 1.1], in which Sun and Cao obtain a zero-one criterion (without asymptotic information) for the shrinking target set associated to the multibase expansion with a fixed $x_{0}$.
@@ 2.6. Proof of Theorem C
Theorem C will be deduced as an application of Theorem B.
Proof of Theorem C.
@@
We now illustrate the link between the shrinking target set and the digit set $D(\xi)$. Given a point $x=[d_{1},\dots,d_{n},\dots]\in[0,1)$, the map $T_{n}$ acts as follows on $x$:
$T_{n}(x)=\frac{d_{n+1}}{b_{n+1}}\ +\ \frac{d_{n+2}}{b_{n+1}b_{n+2}}\ +\ \dots\ +\ \frac{d_{n+\xi(n)}}{b_{n+1}\dots b_{n+\xi(n)}}\ +\ \frac{d_{n+1+\xi(n)}}{b_{n+1}\dots b_{n+1+\xi(n)}}\ +\ \dots\,$
To force $d_{n+1}=d_{n+2}=\cdots=d_{n+\xi(n)}=0$, it suffices to take
$|T_{n}(x)|<r_{n}\quad\text{where}\quad r_{n}:=\frac{1}{b_{n+1}\cdots b_{n+\xi(n)+1}}\,.$
This inequality is equivalent to the statement that
$\quad T_{n}(x)\ \in\ B\left(0,r_{n}\right)\,.$ (2.10)
For each $n\in{\mathbb{N}}$, by choosing $\xi$ so that
$\frac{b_{n+1}}{b_{n+\xi(n)+2}\dots b_{n+\xi(n+1)+2}}\leq 1\,$, (2.11)
we can ensure that the sequence $\{r_{n}\}_{n\in{\mathbb{N}}}$ is decreasing. Now observe that (2.10) is equivalent to $x\in E_{n}$, where $E_{n}$ is the building block set associated to the shrinking target set $\mathcal{W}(T_{\mathbb{N}},0,\{r_{n}\})$. Thus, it follows that
$x\in\mathcal{W}(T_{\mathbb{N}},0,\{r_{n}\})\implies d_{n+1}=...=d_{n+\xi(n)}=0\text{ \ for infinitely many \ }n\in{\mathbb{N}}\,$
and, in particular, Theorem C follows from a direct application of Theorem B. ∎
As mentioned in the introduction, previous works such as [23] explore the typicality of normal numbers for the multibase expansion. Therefore, the more ‘interesting’ case of Theorem C is when $D(\xi)$ consists of arbitrarily long strings of zeros. We wish to analyse when this case occurs on a set of full measure. If the sequence $\{b_{n}\}_{n\in{\mathbb{N}}}$ is bounded, say $b_{n}\leq K$ for all $n\in{\mathbb{N}}$, then we can always find an increasing, unbounded $\xi$ that satisfies the divergent sum condition. Indeed notice that
$r_{n}=\frac{1}{b_{n+1}...b_{n+1+\xi(n)}}\geq\frac{1}{K^{\xi(n)}}\,.$
Picking $\xi(n)\approx\log_{K}(n)\implies K^{\xi(n)}\approx K^{\log_{K}(n)}=n$, and so we are done. This motivates us to ask the following:
Question 1.
Given a sequence of bases $\{b_{n}\}_{n\in{\mathbb{N}}}$ increasing to infinity, is it possible for the sum $\sum_{n=1}^{\infty}\frac{1}{b_{n+1}\dots b_{n+1+\xi(n)}}$ to diverge if $\xi$ is an increasing and unbounded function?
We first note that if we impose the condition that $\{b_{n}\}_{n\in{\mathbb{N}}}$ is an increasing sequence, then (2.11) is satisfied independently of the choice of $\xi$. The answer then depends on the growth rate of $\{b_{n}\}_{n\in{\mathbb{N}}}$, as we illustrate in the following examples. The first shows that if $\{b_{n}\}_{n\in{\mathbb{N}}}$ is an increasing sequence of power law, then the answer to the question is in the negative.
Example.
Let $b_{n}\approx n^{\epsilon}$ for any $\epsilon>0$. Then we shall show that $\sum_{n=1}^{\infty}r_{n}=\infty$ implies $\xi(n)\leq K$. Suppose, for contradiction, that we could find an unbounded, increasing $\xi(n)$ which made the sum diverge. Then there exists some $n_{0}$ such that for all $n\geq n_{0}$, we have $\xi(n)\geq\lceil 2/\epsilon\rceil$. Hence,
$r_{n}=\frac{1}{b_{n+1}\dots b_{n+1+\xi(n)}}\lesssim\frac{1}{(n^{\epsilon})^{2/\epsilon}}=\frac{1}{n^{2}}\,.$
Clearly, this contradicts the divergence of the sum of $r_{n}$, and hence such a $\xi$ does not exist.
However, if $\{b_{n}\}_{n\in{\mathbb{N}}}$ is an increasing sequence of logarithmic growth, then we can answer the question in the affirmative:
Example.
Let $b_{n}\approx\log(n)$. Then we can choose $\xi(n)\approx\big(\log(n)\big)^{1-\epsilon}$ for some $0<\epsilon<1$ such that $\sum_{n=1}^{\infty}r_{n}=\infty$. Indeed, observe that
$\frac{1}{b_{n+1}\dots b_{n+1+\xi(n)}}\gtrsim\frac{1}{\Bigl(\log(n+(\log(n))^{1-\epsilon}+1)\Bigr)^{(\log(n))^{1-\epsilon}}}\,.$
Notice that
$\Big(\log\big(n+(\log(n))^{1-\epsilon}+1\big)\Big)^{1-\epsilon}=\exp\Big((\log(n))^{1-\epsilon}\log\log\big(n+(\log(n))^{1-\epsilon}+1\big)\Big)\,.$
Then we can find some $n_{0}$ such that for all $n\geq n_{0}$,
$\log\log\big(n+(\log(n))^{1-\epsilon}+1\big)(\log(n))^{1-\epsilon}\leq\frac{1}{2}\log(n)\,$,
and so $r_{n}\geq\frac{1}{\sqrt{n}}$. This proves that the sum diverges as desired.
The above analysis is centred around bases $\{b_{n}\}_{n\in{\mathbb{N}}}$ which are increasing and unbounded. It would be interesting to consider the situation where $\{b_{n}\}_{n\in{\mathbb{N}}}$ is unbounded, but not increasing. As previously mentioned, for single base $b$ expansions, the result of [8, Theorem 5] is similar in flavour to Theorem C. However, their theorem applies to digit blocks that are not necessarily zero. In the context of the proof of Theorem C, this would correspond to allowing the point $x_{0}$ to be arbitrary rather than fixing $x_{0}=0$. As discussed in Section 2.5, the result is more general for the single base expansion because the map $t(x)=bx\mod 1$ acts as a left shift on the digit expansion while leaving the constant base sequence unchanged. Using Theorem B, we can adapt their argument to obtain an asymptotic version of their result, with an explicit error term.